OpenStudy (anonymous):

need an answer please: solve for t: -2(-16t-2)=3t-12

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

t = \[-16/29\]

OpenStudy (anonymous):

do u want to know method

OpenStudy (anonymous):

Multiply = -2(16t-2) = 32t+4=3t-12 subtract 32t=3t-16 29t=-16 divide t=-16/29